You are provided the following information about ABC Ltd:

Number of ordinary shares 100,000

Nominal value per ordinary share 50

Market price per ordinary share 80

Net profit before corporation tax 5,000, 000

Rate of corporation tax 30%

Dividend rate 10%

Required: Calculate



I. Dividend yield



II. Earning per share



III. Dividend cover



IV. P/E ratio (8mks)

b) Discuss the uses of ratios (2mks)
1
Expert's answer
2020-12-01T02:17:23-0500

From the given data:


Total value of shares

= number of ordinary shares × nominal value per share.

= 100,000 shares × $50

= $5,000,000


Annual dividend paid

= Value of shares × dividend rate

= $5,000,000 × 10%

= $500,000


Net profit-after tax

NPAT = Net profit before tax - taxation

= $5,000,000 - 30%

= $5,000,000 × 0.70

= $3,500,000



a) (i) "\bold {Answers}"

"Dividend \space yield =\bold { 6.25\%}"


"\bold {Solutions}"

"Dividend \space Yield = \dfrac {Dividend \space per \space Share (DPS)}{Market \space price \space per \space share}×100\%"


"DPS = \dfrac { Total \space annual \space dividends}{Number \space of \space shares}"


"= \dfrac {\$500,000}{100,000 \space shares}"


"= \$5 \space per \space share"


"\therefore \space Dividend \space yield = \dfrac { \$5/share}{\$80/share}×100\%"


"Dividend \space yield = \bold {6.25\%}"



(ii) "\bold {Answers}"


"Earnings \space per \space share, \\ EPS =\bold { \$35/share }"


"\bold {Solutions}"

"Earnings \space per \space share (EPS) = \dfrac {Net \space profit \space after \space tax}{Number \space of \space shares \space outstanding}"


"EPS = \dfrac {\$3,500,000}{100,000 \space shares}"


"=\bold { \$35 \space per \space share}"




(iii) "\bold {Answers}"


"Dividend \space cover = \bold {7 \space times}"


"\bold {Solutions}"


"Dividend \space cover = \dfrac {Net \space profit \space after \space tax}{Total \space ordinary \space dividends \space paid}"


"Dividends \space cover = \dfrac {\$3,500,000}{\$500,000}"


"= \bold {7 \space times}"



(iv) "\bold {Answers}"


"P/E \space ratio = \bold {2.29 \space times}"


"\bold {Solutions}"

"P/E \space ratio \\= \dfrac {Market \space price \space per \space share}{EPS}"


"= \dfrac {\$80/share}{\$35/share}"


"= 2.28571429 \space times"


"= \bold {2.29 \space times}"



b) Ratios are useful in measuring a firm's performance by providing a historical comparison of the firm's profitability, solvency, liquidity, and efficiency.


Ratios are also useful in making interfirm comparison of performance, which helps a firm to understand its performance relative to competitors in the industry. The ratios can, therefore, be used to improve performance.


Ratios are also used by investors to evaluate the profitability or feasibility of an investment decision by analysing stock market ratios such as dividend yield.